OpenStudy (minecraft40952):

Simplify: -7x^3*8x^2

OpenStudy (anonymous):

you add exponents when you multiply

OpenStudy (jdoe0001):

\(\bf -7x^3\cdot 8x^2\implies -7\cdot 8x^3\cdot x^2\implies -7\cdot 8x^{3+2}\)

OpenStudy (minecraft40952):

i dont get it though

OpenStudy (jdoe0001):

http://www.math-play.com/image-exponents-rules.jpg <--- notice the 3rd rule hmmm ok... so... what part you are not getting?

OpenStudy (anonymous):

multiply the coefficients as normal -7 * 8 = -56

OpenStudy (anonymous):

combine the variables x^3 + x^2 = x^5

OpenStudy (minecraft40952):

it just doesent make sence to me

OpenStudy (jdoe0001):

you'd multiplyi the coefficients and the variables IF they're the same "base" their exponents get added, as the rule shows there

OpenStudy (anonymous):

combine (-56)(x^5)

OpenStudy (minecraft40952):

theres no -56 there or 5

OpenStudy (anonymous):

-7 x 8 = ?

OpenStudy (minecraft40952):

and how do i combine them

OpenStudy (anonymous):

do you know how to multiply?

OpenStudy (minecraft40952):

michael theres exponents there

OpenStudy (minecraft40952):

yes

OpenStudy (anonymous):

multiply the coefficients and add the variables

OpenStudy (minecraft40952):

u cant combine if theres x's and stuff

OpenStudy (anonymous):

you add the x's with the coefficients

OpenStudy (anonymous):

for example 2x^2 * 3x^3 = 6x^5

OpenStudy (minecraft40952):

Still not getting it

OpenStudy (anonymous):

-7x^3 * 8x^2 = (-7*8)*(x^3 + x^2)

OpenStudy (minecraft40952):

What's this?

OpenStudy (anonymous):

(-56)(x^5)

OpenStudy (minecraft40952):

Whats that?

OpenStudy (xxdragoonxx):